# Fine-Tuning a Simulation-Driven Estimator

### v2

@@ III Fine-Tuning the TS Estimator
  • Step 2 — Synthetic data generation. If an OOD is detected, perform a few GN iterations in feature space (regularized toward $\hat{\boldsymbol{\theta}}_{\mathrm{init}}$) to obtain an improved estimate $\hat{\boldsymbol{\theta}}_{\mathrm{GN}}$. Then, derive a confidence region around $\hat{\boldsymbol{\theta}}_{\mathrm{GN}}$; sample an informative cloud of parameters from this region; and simulate the DT at these samples to form a small synthetic dataset of feature–parameter pairs.
@@
III-A Step 1: Feature–Space OOD Detector
@@
$\begin{split}\boldsymbol{\mu}_{h}&:=\frac{1}{K}\sum_{k=1}^{K}\boldsymbol{H}_{k:}^{\top},\\ \boldsymbol{S}_{h}&:=\frac{1}{K-1}\sum_{k=1}^{K}\big[\boldsymbol{x}^{(k)}-\boldsymbol{\mu}_{h}\big]\big[\boldsymbol{x}^{(k)}-\boldsymbol{\mu}_{h}\big]^{\top}\\ &\quad\qquad\qquad\qquad+\varepsilon\boldsymbol{I}_{n},\;\varepsilon>0,\end{split}$
where $\boldsymbol{H}_{k:}=(\boldsymbol{x}^{(k)})^{\top}$, and $\boldsymbol{I}_{n}$ is an $n\times n$ identity matrix. The covariance matrix is then used to define a whitening transform, $\boldsymbol{W}:=\boldsymbol{S}_{h}^{-1/2}$, which normalizes the feature space by accounting for correlations and different scales among the features. The test statistic for our observed data is the whitened squared residual norm:
${s}_{\mathrm{obs}}=\|\boldsymbol{W}(\boldsymbol{x}_{\mathrm{obs}}-\boldsymbol{\mu}_{h})\|_{2}^{2}.$ (10)
@@
$s_{k}=\|\boldsymbol{W}(\boldsymbol{H}_{k:}^{\top}-\boldsymbol{\mu}_{h})\|_{2}^{2},\quad k=1,\dots,K.$ (11)
@@
III-B Step 2: Synthetic data generation
If $s_{\mathrm{obs}}>q_{1-\alpha}$, then the compressed features from the first-stage of TS evaluated at $\hat{\boldsymbol{\theta}}_{\mathrm{init}}$ fail to explain $\boldsymbol{x}_{\mathrm{obs}}$ in the whitened space. In that case, we refine $\hat{\boldsymbol{\theta}}_{\mathrm{init}}$ by running Gauss–Newton (GN) iterations that minimize the whitened feature residual between the observed feature and the first-stage–generated feature. Because this residual is random, the Gauss–Newton estimate, denoted $\hat{\boldsymbol{\theta}}_{\mathrm{GN}}$, still carries parameter uncertainty. To effectively fine-tune the second stage of the TS estimator, we need to generate a cloud of synthetic data that covers the region of plausible parameter values around this estimate. We shall now describe precisely the synthetic dataset generation for fine-tuning the second stage of TS as follows.
Gauss–Newton update
We compute the GN update starting from the TS initial estimate. Define a seed-averaged simulator feature map over a small seed set $\mathcal{S}_{\mathrm{avg}}$:
@@ III-B Step 2: Synthetic data generation
$\boldsymbol{r}(\boldsymbol{\theta})$ $=\boldsymbol{W}\big(\bar{\boldsymbol{f}}(\boldsymbol{\theta})-\boldsymbol{x}_{\mathrm{obs}}\big)\in\mathbb{R}^{n},$
where $\boldsymbol{z}(\boldsymbol{\theta},\omega;\boldsymbol{u})$ is the input-output trajectory generated by the simulator for a specific parameter vector $\boldsymbol{\theta}$, a random seed $\omega$, and fixed inputs $\boldsymbol{u}$. Let $\boldsymbol{J}(\boldsymbol{\theta}):=\partial\bar{\boldsymbol{f}}(\boldsymbol{\theta})/\partial\boldsymbol{\theta}\in\mathbb{R}^{n\times d}$ and $\boldsymbol{J}_{w}(\boldsymbol{\theta}):=\boldsymbol{W}\,\boldsymbol{J}(\boldsymbol{\theta})$. A first-order Taylor expansion gives $\boldsymbol{r}(\boldsymbol{\theta}+\boldsymbol{\delta})\approx\boldsymbol{r}(\boldsymbol{\theta})+\boldsymbol{J}_{w}(\boldsymbol{\theta})\,\boldsymbol{\delta}$. We minimize the regularized objective
@@
whose quadratic approximation yields the normal equations
$\underbrace{\big(\boldsymbol{J}_{w}^{\top}\boldsymbol{J}_{w}+\gamma\boldsymbol{I}_{d}\big)}_{=:\boldsymbol{G}_{\gamma}}\,\boldsymbol{\delta}=-\big(\boldsymbol{J}_{w}^{\top}\boldsymbol{r}(\boldsymbol{\theta})+\gamma(\boldsymbol{\theta}-\hat{\boldsymbol{\theta}}_{\mathrm{init}})\big),$ (14)
and the regularized GN step
$\boldsymbol{\delta}^{*}=-\,\boldsymbol{G}_{\gamma}^{-1}\,\big(\boldsymbol{J}_{w}^{\top}\boldsymbol{r}(\boldsymbol{\theta})+\gamma(\boldsymbol{\theta}-\hat{\boldsymbol{\theta}}_{\mathrm{init}})\big),$ (15)
followed by a backtracking line search (e.g., Armijo). A few iterations yield $\hat{\boldsymbol{\theta}}_{\mathrm{GN}}$.
@@
where $\chi^{2}_{d}(\beta)$ is the chi-squared quantile and $\boldsymbol{G}\approx\boldsymbol{J}_{w}^{\top}\boldsymbol{J}_{w}+\lambda\boldsymbol{I}_{d}$ is the approximate Fisher Information Matrix (FIM). This ellipsoid represents a high-confidence region, and sampling from it is statistically efficient for exploring well-identified parameters. However, this primary method can fail if the model is insensitive to one or more parameters. In such cases, the FIM, $\boldsymbol{G}$, becomes ill-conditioned, causing the confidence ellipsoid to be stretched along the insensitive directions, and thereby making the sampling procedure numerically unstable. To overcome this, we alternate between this sampling approach and sensitivity sampling. The latter sampling procedure handles the ill-conditioned directions. It works by generating samples along individual parameter axis with a step size that is inversely proportional to the simulator’s sensitivity to that parameter. The sensitivity of coordinate $j$ is defined as $s^{\prime}_{j}:=\|\boldsymbol{J}_{w}\boldsymbol{e}_{j}\|_{2}$. To set the sampling scale, we define two quantities. First, $\Delta_{\mathrm{misfit}}:=\|\boldsymbol{r}(\hat{\boldsymbol{\theta}}_{\mathrm{GN}})\|_{2}$, measures the discrepancy between the observed summary statistic and the one generated by the simulator at the point estimate $\hat{\boldsymbol{\theta}}_{\mathrm{GN}}$. Second, we compute $d_{\mathrm{med}}:=\mathrm{median}_{k}\,d_{k}$, based on the spread of pre-computed summary features, where each $d_{k}:=\|\boldsymbol{W}(\boldsymbol{H}_{k:}-\boldsymbol{\mu}_{h})\|_{2}$ represents the whitened residual. We then set coordinate scales, $r_{j}$, that increase exploration for less sensitive parameters (small $s^{\prime}_{j}$) while also adapting to the simulator’s noise level and the current feature mismatch as
$r_{j}\;=\;\frac{d_{\mathrm{med}}}{s^{\prime}_{j}}\,\Delta_{\mathrm{misfit}}.$ (17)
Our final sampling procedure combines confidence ellipsoid and sensitivity sampling, yielding a hybrid sampling scheme, given by $\bar{\boldsymbol{\theta}}_{k}=\hat{\boldsymbol{\theta}}_{\mathrm{GN}}+\boldsymbol{\Delta\theta}_{k}$, where $\boldsymbol{\Delta\theta}_{k}$ is drawn as
$\text{(confidence set sampling)}\quad\boldsymbol{\Delta\theta}_{k}=\boldsymbol{L}\,\boldsymbol{z},\;\;\boldsymbol{z}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{d}),\;\;$
$\quad\quad\qquad\qquad\qquad\qquad\qquad\qquad\quad\boldsymbol{L}\boldsymbol{L}^{\top}\approx\boldsymbol{G}^{-1},$
@@
where $\boldsymbol{e}_{i}\in\mathbb{R}^{d}$ is the $i$-th standard basis vector ($1$ at position $i$, zeros elsewhere), and $\boldsymbol{L}$ is the Cholesky factor of $\boldsymbol{G}^{-1}$. To create the fine-tuning dataset, $\mathcal{D}_{\mathrm{fine}}$, we generate a synthetic observation $\bar{\boldsymbol{z}}_{k}$ from the simulator for each parameter sample $\bar{\boldsymbol{\theta}}_{k}$. Each observation is then compressed using $h(\cdot)$, forming a dataset of pairs: $\mathcal{D}_{\mathrm{fine}}=\{(h(\boldsymbol{\bar{z}}_{k}),\bar{\boldsymbol{\theta}}_{k})\}_{k=1}^{M_{\mathrm{FT}}}$.
Remark III.1
The quantity $\Delta_{\mathrm{misfit}}$ can be artificially small due to noise induced by the randomness of the simulator $\mathcal{M}$, so the naive radius $\Delta_{\mathrm{misfit}}/s^{\prime}_{j}$ would under–explore. Multiplying by $d_{\mathrm{med}}:=\mathrm{median}_{k}\|\boldsymbol{W}(\boldsymbol{H}_{k:}-\boldsymbol{\mu}_{h})\|_{2}$ rescales the step to the typical feature fluctuation at $\hat{\boldsymbol{\theta}}_{\mathrm{init}}$, giving $r_{j}=(d_{\mathrm{med}}/s^{\prime}_{j})\,\Delta_{\mathrm{misfit}}$. This preserves the desired inverse–sensitivity ($1/s^{\prime}_{j}$) while making the samples noise–aware.
III-C Step 3: Final–Layer Transfer on a Targeted Set
@@
$\begin{split}\min_{\boldsymbol{\Psi}}\;\frac{1}{M_{\mathrm{FT}}}\sum_{m=1}^{M_{\mathrm{FT}}}&\sum_{j=1}^{d}\tilde{w}_{j}\,L\Big(\big[g_{\mathrm{pre}}(h(\bar{\boldsymbol{z}}_{m});\boldsymbol{\phi},\boldsymbol{\Psi})\big]_{j},\ [\bar{\boldsymbol{\theta}}_{m}]_{j}\Big)\;\\ &+\lambda_{\mathrm{orth}}\!\!\sum_{1\leq j<k\leq d}\!\!\big\|\boldsymbol{U}_{j,1}\boldsymbol{U}_{k,1}^{\top}\big\|_{F}^{2}.\end{split}$ (19)
IV Simulation Study
We evaluate the proposed fine-tuning scheme on two nonlinear systems: (i) the Van der Pol (VDP) oscillator and (ii) cascaded water tanks. In both cases, the model structure is known but the true parameters may lie outside the sampling set used to pretrain the TS estimator. We compare against TS-pre (pretrained TS), the Prediction Error Method (PEM) [11], and a dual Extended Kalman Filter [17] (dual-EKF). For PEM and dual-EKF we report results under three initializations: Good Init (a Gaussian perturbation of the true parameter value), TS-Init (initialized at $\hat{\theta}_{\mathrm{pre}}(\boldsymbol{z}_{0})$), and Bad/Worse-Case Init (point far from $\boldsymbol{\theta}_{0}$).
@@ IV Simulation Study
For all simulations we set $K=150$ in (III-A). The whitened Jacobian $\boldsymbol{J}_{w}$ is obtained via finite differences. To compute $\bar{\boldsymbol{f}}(\boldsymbol{\theta})$ in (12) and the Jacobians, we use common random numbers with the seed set $S_{\mathrm{avg}}=\{8000,8001,8002\}$. Both examples are simulated in discrete time using forward Euler with a fixed step size $\Delta t$ (VDP: $0.05$ s; tanks: $0.5$ s). The Gauss–Newton regularization parameter in (13) is $\gamma=10^{-3}$ (VDP) and $\gamma=10^{-4}$ (tanks). The orthogonality regularizer $\lambda_{\mathrm{orth}}$ on the first head layer is 0 for VDP (single output) and $5\times 10^{-4}$ for the tanks.
@@
For VDP we use $\mathrm{AR}(5)$ coefficients fitted to $y$; for the tanks we use $\mathrm{ARX}(32,32)$ coefficients fitted to $(y,u)$. The stage-2 regressor $g_{\mathrm{pre}}$ is a fully connected network with a shared trunk and one head per parameter, and the following settings:
@@
IV-A Van der Pol Oscillator

### v3

@@ 4 Fine-Tuning the TS Estimator
  • Step 2 — Synthetic data generation. If an OOD is detected, perform a few Gauss-Newton (GN) iterations in feature space, regularized toward $\hat{\boldsymbol{\theta}}_{\mathrm{init}}$, to obtain an improved estimate $\hat{\boldsymbol{\theta}}_{\mathrm{GN}}$. Then, derive a confidence region around $\hat{\boldsymbol{\theta}}_{\mathrm{GN}}$; sample an informative cloud of parameters from this region; and simulate the DT at these samples to form a small synthetic dataset of feature–parameter pairs.
@@
4.1 Step 1: Feature–Space OOD Detector
@@
$\begin{split}\boldsymbol{\mu}&:=\frac{1}{K}\sum_{k=1}^{K}\boldsymbol{H}_{k:}^{\top},\\ \boldsymbol{S}&:=\frac{1}{K-1}\sum_{k=1}^{K}\big[\boldsymbol{x}^{(k)}-\boldsymbol{\mu}\big]\big[\boldsymbol{x}^{(k)}-\boldsymbol{\mu}\big]^{\top}\\ &\qquad\qquad\qquad\qquad+\varepsilon\boldsymbol{I}_{n},\;\varepsilon>0,\end{split}$ (8)
where $\boldsymbol{H}_{k:}=(\boldsymbol{x}^{(k)})^{\top}$, and $\boldsymbol{I}_{n}$ is an $n\times n$ identity matrix. The covariance matrix is then used to define a whitening transform, $\boldsymbol{S}^{-1/2}$, which normalizes the feature space by accounting for correlations and different scales among the features. The test statistic for our observed data is the whitened squared residual norm:
$s_{\mathrm{obs}}=\big\|\boldsymbol{S}^{-1/2}\big(\boldsymbol{x}_{\mathrm{obs}}-\boldsymbol{\mu}\big)\big\|_{2}^{2},$ (9)
@@
$s_{k}=\big\|\boldsymbol{S}^{-1/2}\big(\boldsymbol{H}_{k:}^{\top}-\boldsymbol{\mu}\big)\big\|_{2}^{2}.$ (10)
@@
4.2 Step 2: Synthetic data generation
If $s_{\mathrm{obs}}>q_{1-\alpha}$, then the compressed features from the first-stage of TS evaluated at $\hat{\boldsymbol{\theta}}_{\mathrm{init}}$ fail to explain $\boldsymbol{x}_{\mathrm{obs}}$ in the whitened space. In that case, we refine $\hat{\boldsymbol{\theta}}_{\mathrm{init}}$ by running GN iterations that minimize the whitened feature residual between the observed feature and the first-stage generated feature. Because this residual is random, the GN estimate, denoted $\hat{\boldsymbol{\theta}}_{\mathrm{GN}}$, still carries parameter uncertainty. To effectively fine-tune the second stage of the TS estimator, we need to generate a cloud of synthetic data that covers the region of plausible parameter values around this estimate. We shall now describe precisely the synthetic dataset generation for fine-tuning the second stage of TS as follows.
GN update
We compute the GN update starting from the TS initial estimate $\hat{\boldsymbol{\theta}}_{\mathrm{init}}$. We define a seed-averaged simulator feature map $\bar{\boldsymbol{f}}(\boldsymbol{\theta})$ and its corresponding whitened residual $\boldsymbol{r}(\boldsymbol{\theta})$:
@@ 4.2 Step 2: Synthetic data generation
$\boldsymbol{r}(\boldsymbol{\theta})$ $=\boldsymbol{S}^{-1/2}\big(\bar{\boldsymbol{f}}(\boldsymbol{\theta})-\boldsymbol{x}_{\mathrm{obs}}\big)\in\mathbb{R}^{n},$
where $\boldsymbol{z}(\boldsymbol{\theta},\omega;\boldsymbol{u})$ is the simulator trajectory for parameter $\boldsymbol{\theta}$ and seed $\omega$. We then minimize the following regularized objective function using iterative GN steps (with a backtracking line search):
@@
A few iterations of minimizing this objective yield the refined estimate $\hat{\boldsymbol{\theta}}_{\mathrm{GN}}$. Let $\partial\bar{\boldsymbol{f}}(\boldsymbol{\theta})/\partial\boldsymbol{\theta}\in\mathbb{R}^{n\times d}$ denote the Jacobian of the seed-averaged feature map, and let $\boldsymbol{S}(\boldsymbol{\theta})$ be the feature covariance estimated from simulator draws at parameter $\boldsymbol{\theta}$, analogous to Eq. (8). Define the whitened Jacobian  $\boldsymbol{J}\;\coloneqq\;\boldsymbol{S}^{-1/2}\,\partial\bar{\boldsymbol{f}}(\boldsymbol{\theta})/\partial\boldsymbol{\theta}.$ We re-compute $\boldsymbol{\mu},\boldsymbol{S}$ and $\boldsymbol{J}$ at $\hat{\boldsymbol{\theta}}_{\mathrm{GN}}$, and set $\boldsymbol{G}\coloneqq\boldsymbol{J}^{\top}\boldsymbol{J}+\lambda\boldsymbol{I}_{d}.$
@@
where $\chi^{2}_{d}(\beta)$ is the chi-squared quantile and $\boldsymbol{G}$ is the approximate Fisher Information Matrix (FIM). This ellipsoid represents a high-confidence region, and sampling from it is statistically efficient for exploring well-identified parameters. However, this primary method can fail if the model is insensitive to one or more parameters. In such cases, the FIM, $\boldsymbol{G}$, becomes ill-conditioned, causing the confidence ellipsoid to be stretched along the insensitive directions, and thereby making the sampling procedure numerically unstable. To overcome this, we alternate between this sampling approach and sensitivity sampling. The latter sampling procedure handles the ill-conditioned directions. It works by generating samples along individual parameter axis with a step size that is inversely proportional to the simulator’s sensitivity to that parameter. The sensitivity of coordinate $j$ is defined as $s^{\prime}_{j}:=\|\boldsymbol{J}\boldsymbol{e}_{j}\|_{2}$. To set the sampling scale, we define two quantities. First, $\Delta_{\mathrm{misfit}}:=\|\boldsymbol{r}(\hat{\boldsymbol{\theta}}_{\mathrm{GN}})\|_{2}$, measures the discrepancy between the observed summary statistic and the one generated by the simulator at the point estimate $\hat{\boldsymbol{\theta}}_{\mathrm{GN}}$. Second, we compute $d_{\mathrm{med}}:=\mathrm{median}_{k}\,d_{k}$, based on the spread of pre-computed summary features, where each $d_{k}:=\|\boldsymbol{S}^{-1/2}(\boldsymbol{H}_{k:}-\boldsymbol{\mu})\|_{2}$ represents the whitened residual. We then set coordinate scales, $r_{j}$, that increase exploration for less sensitive parameters (small $s^{\prime}_{j}$) while also adapting to the simulator’s noise level and the current feature mismatch as $r_{j}\;=\;\frac{d_{\mathrm{med}}}{s^{\prime}_{j}}\,\Delta_{\mathrm{misfit}}.$
Our sampling procedure combines confidence ellipsoid and sensitivity sampling, yielding a hybrid sampling scheme, given by $\bar{\boldsymbol{\theta}}_{k}=\hat{\boldsymbol{\theta}}_{\mathrm{GN}}+\boldsymbol{\Delta\theta}_{k}$, where $\boldsymbol{\Delta\theta}_{k}$ is drawn as
$\text{(confidence set sampling)}\quad\boldsymbol{\Delta\theta}_{k}=\boldsymbol{B}\,\boldsymbol{z},\;\;\boldsymbol{z}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{d}),\;\;$
$\quad\quad\qquad\qquad\qquad\qquad\qquad\qquad\quad\boldsymbol{B}\boldsymbol{B}^{\top}\approx\boldsymbol{G}^{-1},$
@@
where $\boldsymbol{e}_{i}\in\mathbb{R}^{d}$ is the $i$-th standard basis vector ($1$ at position $i$, zeros elsewhere), and $\boldsymbol{B}$ is the Cholesky factor of $\boldsymbol{G}^{-1}$. To create the fine-tuning dataset, $\mathcal{D}_{\mathrm{fine}}$, we generate a synthetic observation $\bar{\boldsymbol{z}}_{k}$ from the simulator for each parameter sample $\bar{\boldsymbol{\theta}}_{k}$. Each observation is then compressed using $h$, forming a dataset of pairs: $\mathcal{D}_{\mathrm{fine}}=\{(h(\boldsymbol{\bar{z}}_{k}),\bar{\boldsymbol{\theta}}_{k})\}_{k=1}^{M_{\mathrm{FT}}}$, where $M_{\mathrm{FT}}$ is the number of new synthetic samples.
Remark 4.1
The quantity $\Delta_{\mathrm{misfit}}$ can be artificially small due to noise induced by the randomness of the simulator $\mathcal{M}$, so the naive radius $\Delta_{\mathrm{misfit}}/s^{\prime}_{j}$ would under–explore. Multiplying by $d_{\mathrm{med}}$ rescales the step to the typical feature fluctuation at $\hat{\boldsymbol{\theta}}_{\mathrm{init}}$, giving $r_{j}=(d_{\mathrm{med}}/s^{\prime}_{j})\,\Delta_{\mathrm{misfit}}$. This preserves the desired inverse–sensitivity ($1/s^{\prime}_{j}$) while making the samples noise–aware.
Remark 4.2 (Theoretical justification of Steps 1 and 2)
For Step 1, under the null hypothesis (in-distribution), the observed features $\boldsymbol{x}_{\mathrm{obs}}$ and the simulated features $\boldsymbol{x}^{(k)}$ are asymptotically normal with the same mean and covariance. This implies that their whitened residuals ($s_{\mathrm{obs}}$ and $s_{k}$) converge to the same $\chi^{2}$ distribution, validating the bootstrap test. For Step 2, the GN objective $v(\boldsymbol{\theta})$ is an M-estimator whose expectation is minimized at the true $\boldsymbol{\theta}_{0}$. Therefore, the GN estimate $\hat{\boldsymbol{\theta}}_{\mathrm{GN}}$ is consistent [16, Ch. 5], and the FIM $\boldsymbol{G}$, as a consistent estimator of the asymptotic covariance, defines an asymptotically valid confidence ellipsoid for sampling.
4.3 Step 3: Final–Layer Transfer on a Targeted Set
@@
$\hat{\boldsymbol{\Psi}}\in\arg\min_{\boldsymbol{\Psi}}\;\frac{1}{M_{\mathrm{FT}}}\sum_{i=1}^{M_{\mathrm{FT}}}$ $\sum_{j=1}^{d}\tilde{w}_{j}\ell([g_{\mathrm{pre}}(h(\bar{\boldsymbol{z}}_{i});\boldsymbol{\phi},\boldsymbol{\Psi})]_{j},[\bar{\boldsymbol{\theta}}_{i}]_{j})$ (15)
$+\lambda_{\mathrm{orth}}\!\!\sum_{1\leq j<k\leq d}\!\!\big\|\boldsymbol{U}_{j,1}\boldsymbol{U}_{k,1}^{\top}\big\|_{F}^{2},$
where $\tilde{w}_{j}\;=\;d\,[\boldsymbol{G}]_{jj}/\sum_{k=1}^{d}[\boldsymbol{G}]_{kk}$, with $\boldsymbol{G}]_{jj}$ denoting the $j$-th diagonal entry of $\boldsymbol{G}$, and $\ell$ is the Huber loss function.
5 Simulation Study
We evaluate the proposed fine-tuning scheme on two nonlinear systems: (i) the Van der Pol (VDP) oscillator and (ii) cascaded water tanks. In both cases, the model structure is known but the true parameters may lie outside the sampling set used to pretrain the TS estimator. We compare the fine-tuned TS (TS-fine) against pretrained TS (TS-pre), the PEM, and a dual-EKF. For PEM and dual-EKF we report results under three initializations: Good Init (a Gaussian perturbation of the true parameter value), TS-Init (initialized at $\hat{f}_{\mathrm{pre}}(\boldsymbol{z}_{0})$), and Bad/Worse-Case Init (point far from $\boldsymbol{\theta}_{0}$).
@@ 5 Simulation Study
For all simulations we set $K=150$ in (8). The whitened Jacobian $\boldsymbol{J}$ is obtained via finite differences. To compute $\bar{\boldsymbol{f}}(\boldsymbol{\theta})$ in (11) and the Jacobians, we use common random numbers with the seed set $S_{\mathrm{avg}}=\{8000,8001,8002\}$. Both examples are simulated in discrete time using forward Euler with a fixed step size $\Delta t$ (VDP: $0.05$ s; tanks: $0.5$ s). The Gauss–Newton regularization parameter in (12) is $\gamma=10^{-3}$ (VDP) and $\gamma=10^{-4}$ (tanks). The orthogonality regularizer $\lambda_{\mathrm{orth}}$ on the first head layer is 0 for VDP (single output) and $5\times 10^{-4}$ for the tanks.
@@
For VDP we use $\mathrm{AR}(5)$ coefficients fitted to $y$; for the tanks we use $\mathrm{ARX}(64,64)$ coefficients fitted to $(y,u)$. The stage-2 regressor $g_{\mathrm{pre}}$ is a fully connected network with a shared trunk and one head per parameter, and the following settings:
@@
Remark 5.1
We fix all hyperparameters ($K$, $M_{\mathrm{FT}}$, $\gamma$, $\lambda_{\mathrm{orth}}$, architectures, and training choices) a priori across all experiments.
5.1 Van der Pol Oscillator